# Negative Pre-aware for Noisy Cross-modal Matching

### 10 Dec 2023 v1

@@ Abstract
Cross-modal noise-robust learning is a challenging task since noisy correspondence is hard to recognize and rectify. Due to the cumulative and unavoidable negative impact of unresolved noise, existing methods cannot maintain a stable performance when the noise increases. In this paper, we present a novel Negative Pre-aware Cross-modal (NPC) matching solution for large visual-language model fine-tuning on noisy downstream tasks. It is featured in two aspects: (1) For noise recognition and resistance, previous methods usually directly filter out a noise subset, we propose to estimate the negative impact of each sample. It does not need additional correction mechanisms that may predict unreliable correction results, leading to self-reinforcing error. We assign a confidence weight to each sample according to its negative impact in the training process. This adaptively adjusts the contribution of each sample to avoid noisy accumulation. (2) For maintaining stable performance with increasing noise, we utilize the memorization effect of DNNs by maintaining a memory bank. Specifically, we apply GMM to select high-confident clean samples as the memory entry, where the memory entry is used to estimate the negative impact of each sample. Since clean samples are easier distinguished by GMM with increasing noise, the memory bank can still maintain high quality at a high noise ratio. Compared to the correction mechanism focusing on noise samples, memory bank-based estimation is more robust, which makes the model performance stable on noisy datasets. Extensive experiments demonstrate that our method significantly improves matching accuracy and performance stability at increasing noise ratio. Our approach also surpasses the state-of-the-art methods by a large margin.
@@ Introduction
These works have achieved promising performance by training on large-scale datasets. However, it is expensive to obtain a well-annotated dataset in practical scenarios. The manual-annotated datasets, e.g., MSCOCO (Lin et al. 2014) and Flickr30K (Young et al. 2014), incorporate a significant number of inaccurate descriptions, namely noise correspondence. Unlike noise label in classification tasks, the noise here is mismatched cross-modal pairs which is more difficult to deal with, since involves both visual and textual modeling. Therefore, a series of approaches (Huang et al. 2021; Yang et al. 2023; Han et al. 2023) following the noise-rectify paradigm have been developed to counter the negative impact of the noise. These methods typically filter out the noise subset from the original training set. Subsequently, they address the noise issue through label correction (Huang et al. 2021; Yang et al. 2023; Han et al. 2023). Nevertheless, the inherent flaw of the noise-rectify paradigm cannot maintain the performance stability in the existence of severe noise. As shown in Fig. 1(b), we compare the performance of different methods using R@1 metric, including noise-rectify based approaches (Huang et al. 2021; Yang et al. 2023), the large-scale CLIP-based approaches (Radford et al. 2021; Chun 2023) and our approach. We employ variance ($var$) of R@1 at different noise ratio to illustrate the “performance stability”. Obviously, noise-rectify based methods exhibit unstable performance with a considerably larger variance than ours. Additionally, CLIP-based methods also lack consistent performance with increasing noise, even though CLIP is a powerful pre-trained model. Most existing noise-rectify paradigms rely on collaborative rectifying with multiple models. Since the limitation of the rectifying mechanism, the matching performance under high-noise is unstable. In these works (Huang et al. 2021; Yang et al. 2023), the new labels are entirely estimated by DNN models. With high noise ratio, some indistinguishable noise correspondences are prone to be directly learned and remembered by DNNs, ultimately leading to a dramatic drop in performance under high-noise. Existing methods emphasize the “discrimination learning” ability but ignore the “stability”. In our opinion, two essential abilities are required for noise-robust learning for large visual-language model fine-tuning on noisy downstream tasks: 1) the ability to distinguish between noisy and clean samples, 2) maintain the stability of “discriminative learning” with increasing noise.
@@
To address aforementioned challenges, we propose a novel approach named Negative Pre-aware Cross-modal (NPC) matching. NPC adopts a unique Negative Pre-aware (NP) paradigm for robust learning. Unlike previous paradigms that mainly focus on noise filtering or correction, the NP paradigm adaptively assesses the potential negative impact of each sample before the model learning (see Fig. 1(a)). DNNs tends to prioritize learning easy samples over noisy and challenging ones (Arpit et al. 2017; Xia et al. 2021). With gradually fitting noise samples, the model begins to generate incorrect predictions (Liu et al. 2020).
In other words, once the model learned a noise pair, fitting certain specific clean samples becomes more challenging. These clean samples usually have images or texts that are similar to the noise pair.
@@
Inspired by this phenomenon, our NPC uses easy-distinguishable clean samples to estimate negative impacts. We rigorously choose a reliable clean subset from the training data by using Gaussian Mixture Model (Li, Socher, and Hoi 2020; Permuter, Francos, and Jermyn 2006) to fit the loss distribution of each pair. Samples with a clean probability close to 1 are considered clean, where high-confident clean samples are maintained in a gradually updated Memory Bank ($MB$). Thus the $MB$ could be used to assist the model in estimating negative impact before to fully model training. Correspondingly, a small confidence weight will be assigned to high-negative samples. Extensive experiments on large-scale pre-trained CLIP have verified its effectiveness under various settings.
@@
  • Extensive experiments are conducted on two challenging datasets, showcasing the NPC’s superiority over the state-of-the-art methods. Moreover, with the increasing noise, both quantitative and qualitative results affirm that NPC demonstrates notably higher performance stability compared to previous methods.
@@
Typical image-text matching methods align data from different modalities to measure similarity. Early works (Faghri et al. 2018; Song and Soleymani 2019; Wang et al. 2018; Qian et al. 2021) mainly focus on global alignments. VSE++ (Faghri et al. 2018) uses relative hard pairs to enhance encoders. PIE-Nets (Song and Soleymani 2019) generate diverse representations for robust embeddings. Some prior works adopt attention mechanisms to achieve fine-grained local alignments. SCAN (Lee et al. 2018) pioneers cross-attention for fine-grained modality alignment. A line of SCAN-based methods (Li et al. 2019a; Chen et al. 2020; Liu et al. 2019; Diao et al. 2021; Zhang et al. 2023) further enhance the fine-grained matching accuracy. Subsequently, many works (Chun et al. 2021; Chun 2023; Li et al. 2022) devote to modeling the many-to-many relationships in image-caption pairs.
Recently, with the success of transformer-based language models (Devlin et al. 2019) and vision transformer (Dosovitskiy et al. 2021), vision-language pre-training (VLP) models, such as CLIP (Radford et al. 2021), have shown strong performance in multiple cross-modal tasks (Jiang and Ye 2023; Gao et al. 2022; You et al. 2023). Although VLP possesses impressive zero-shot ability, it still reveals vulnerabilities in training with noisy datasets on specific downstream tasks. In this paper, we employ CLIP as our backbone and introduce an anti-noise learning strategy.
@@
Huang et al. (Huang et al. 2021) first tackle noise correspondences, which consider mismatched cross-modal pairs instead of incorrect annotations. Since then, several approaches following the noise-rectify process, which can be categorized into two groups: noise correction and noise re-weighting. Noise correction methods achieve robust learning by correcting the similarity (Han et al. 2023) or correspondence label (Huang et al. 2021; Yang et al. 2023) of noise pairs. The noise re-weighting methods (Qin et al. 2022) degrade the contribution of noise samples to achieve robust learning. All these methods require splitting a noise subset from the original training dataset. Subsequently, they proceed with rectification within this subset. Nonetheless, as noise increases, the imprecise subset division and inaccurate rectification can amplify adverse effects. Different from these works, we sidestep the problem by forecasting per-sample negative impact following the novel NP paradigm.
@@
Fig. 1 provides a clear visual representation, demonstrating that when the noise ratio rises from 20% to 60%, the CLIP’s R@1 performance experiences a steep decline from 82.3% to 66.3%. Therefore, the NPC approach is introduced to enhance the stability and robustness of pre-trained models in tackling noise challenges. The training pipeline, depicted in Fig. 2, comprises the two main components that will be elaborated upon in the subsequent section. Furthermore, the pseudo-code of the Negative Pre-aware Cross-Modal training process is shown in Alg. 1.
Input: The noisy training dataset $D=\{(I_{i},T_{i})\}_{i=1}^{N}$, the batch size $m$, the base model $A$.
1 Memory Bank Construction:
2 select $D_{c}$ from $\text{GMM}(D,A)$
3 for $i\in\{1,\dots,N\}$ do
4      $MB[i]=((I_{i}^{I},T_{i}^{I}),(I_{i}^{T},T_{i}^{T}))$;
5 end for
6for $s\in\{1,\dots,N/m\}$ do
7        // Sample a $batch$ and $MB_{batch}$.
8       $D_{batch}=\{(I_{j},T_{j})\}_{j=1}^{m}$,$MB_{batch}=\{b_{j}\}_{j=1}^{m}$;
9       Negative Impact-aware:
10       $A^{\prime}_{s}=A_{s}\tt{.clone()}$;
11       compute loss $p^{s},q^{s}$ on $MB_{batch}$;
12       $A^{\prime}_{s+1}=A^{\prime}_{s}\tt{.update()}$, constrained by $\mathcal{L}_{CE}$;
13       compute loss $p^{s+1},q^{s+1}$ on $MB_{batch}$;
14       $r\leftarrow(p^{s},~{}q^{s},~{}p^{s+1},~{}q^{s+1})$ by Eq. 7;
15       for $r_{k}\in r$ do
16            if $r_{k}<1$ then $w_{k}=\tanh(r_{k})$;
17            else $w_{k}=1$;
18
19       end for
20      Re-training:
21       $\mathcal{L}_{total}\leftarrow w\mathcal{L}_{CE}+\mathcal{L}_{MB}$;
22       $A_{s+1}=A_{s}\tt{.update()}$, constrained by $\mathcal{L}_{total}$;
23
24 end for
Output: Image-text matching model $A$.
Algorithm 1 Negative Pre-aware Cross-modal (NPC)
@@ 21       $\mathcal{L}_{total}\leftarrow w\mathcal{L}_{CE}+\mathcal{L}_{MB}$;
The proposed NPC is evaluated on two benchmark datasets, MSCOCO (Lin et al. 2014) and Flickr30K (Young et al. 2014):
@@
To illustrate the effectiveness of proposed NPC, we compare it with various approaches, including general cross-modal matching methods SCAN (Lee et al. 2018), SAF (Diao et al. 2021), noise-robust learning methods NCR (Huang et al. 2021), DECL (Qin et al. 2022), BiCro (Yang et al. 2023), and CLIP with fine-tuning (Radford et al. 2021). It is worth noting that CLIP is the baseline of our method. The results are shown in Table 1.
It can be seen that NPC significantly outperforms all methods across all noise ratios. Notably, on Flickr30K with 60% noise ratio, NPC outperforms the current state-of-the-art approach BiCro with a large R@1 performance gap. To be specific, the R@1 performance of NPC is 15.4% higher than BiCro on image-text matching (i2t), as well as 16.9% higher than BiCro on text-to-image matching (t2i). Compared to the baseline CLIP, NPC has achieved immense improvement in all metrics and benchmarks. Furthermore, as the noise ratio increases, the performance gap between NPC and baseline becomes larger. For instance, on the MSCOCO 1K set, when the noise ratio ranges from 0% to 60%, the R@1 performance gap between NPC and baseline separately increases from 2.3% to 11.2% on i2t, and 3.3% to 13.4% on t2i. This phenomenon is powerful to prove the effectiveness of NPC on robust learning.

### 15 Dec 2023 v2

@@ Abstract
Cross-modal noise-robust learning is a challenging task since noisy correspondence is hard to recognize and rectify. Due to the cumulative and unavoidable negative impact of unresolved noise, existing methods cannot maintain a stable performance when the noise increases. In this paper, we present a novel Negative Pre-aware Cross-modal (NPC) matching solution for large visual-language model fine-tuning on noisy downstream tasks. It is featured in two aspects: (1) For noise recognition and resistance, previous methods usually directly filter out a noise subset, we propose to estimate the negative impact of each sample. It does not need additional correction mechanisms that may predict unreliable correction results, leading to self-reinforcing error. We assign a confidence weight to each sample according to its negative impact in the training process. This adaptively adjusts the contribution of each sample to avoid noisy accumulation. (2) For maintaining stable performance with increasing noise, we utilize the memorization effect of DNNs by maintaining a memory bank. Specifically, we apply GMM to select high-confident clean samples as the memory entry, where the memory entry is used to estimate the negative impact of each sample. Since clean samples are easier distinguished by GMM with increasing noise, the memory bank can still maintain high quality at a high noise ratio. Compared to the correction mechanism focusing on noise samples, memory bank-based estimation is more robust, which makes the model performance stable on noisy datasets. Extensive experiments demonstrate that our method significantly improves matching accuracy and performance stability at increasing noise ratio. Our approach also surpasses the state-of-the-art methods by a large margin. The code is available at: https://github.com/ZhangXu0963/NPC.
@@ Introduction
These works have achieved promising performance by training on large-scale datasets. However, it is expensive to obtain a well-annotated dataset in practical scenarios. The manual-annotated datasets, e.g., MSCOCO (Lin et al. 2014), Flickr30K (Young et al. 2014), and Conceptual Captions (Sharma et al. 2018), incorporate a significant number of inaccurate descriptions, namely noisy correspondence. Unlike noisy label in classification tasks, the noise here is mismatched cross-modal pairs which is more difficult to deal with, since involves both visual and textual modeling. Therefore, a series of approaches (Huang et al. 2021; Yang et al. 2023; Han et al. 2023) following the noise-rectify paradigm have been developed to counter the negative impact of the noise. These methods typically filter out the noise subset from the original training set, and address the noise issue through label correction. Nevertheless, the inherent flaw of the noise-rectify paradigm cannot maintain the performance stability in the existence of severe noise. As shown in Fig. 1(b), we compare the performance of different methods using R@1 metric, including noise-rectify based approaches (Huang et al. 2021; Yang et al. 2023), the CLIP-based approaches (Radford et al. 2021; Chun 2023) and our approach. We employ variance ($var$) of R@1 at different noise ratio to illustrate the “performance stability”. Obviously, noise-rectify based methods exhibit unstable performance with a considerably larger variance than ours. Additionally, CLIP-based methods also lack consistent performance with increasing noise, even though CLIP is a powerful pre-trained model. Most existing noise-rectify paradigms rely on collaborative rectifying with multiple models. Since the limitation of the rectifying mechanism, the matching performance under high-noise is unstable. In these works (Huang et al. 2021; Yang et al. 2023), the new labels are entirely estimated by DNN models. With high noise ratio, some indistinguishable noise correspondences are prone to be directly learned and remembered by DNNs, ultimately leading to a dramatic drop in performance under high-noise. Existing methods emphasize the “discrimination learning” ability but ignore the “stability”. In our opinion, two essential abilities are required for noise-robust learning for large visual-language model fine-tuning on noisy downstream tasks: 1) the ability to distinguish between noisy and clean samples, 2) maintain the stability of “discriminative learning” with increasing noise.
@@
To address aforementioned challenges, we propose a novel approach named Negative Pre-aware Cross-modal (NPC) matching. NPC adopts a unique Negative Pre-aware (NP) paradigm for robust learning. Unlike previous paradigms that mainly focus on noise filtering or correction, the NP paradigm adaptively assesses the potential negative impact of each sample before the model learning (see Fig. 1(a)). DNNs tends to prioritize learning easy samples over noisy and challenging ones (Arpit et al. 2017; Xia et al. 2021). With gradually fitting noise samples, the model begins to generate incorrect predictions (Liu et al. 2020). In other words, once the model learned a noise pair, fitting certain specific clean samples becomes more challenging. These clean samples usually have images or texts that are similar to the noise pair.
@@
Inspired by this phenomenon, our NPC uses easy-distinguishable clean samples to estimate negative impacts. We rigorously choose a reliable clean subset from the training data by using Gaussian Mixture Model (Li, Socher, and Hoi 2020; Permuter, Francos, and Jermyn 2006) to fit the loss distribution of each pair. And high-confident clean samples are maintained in a Memory Bank ($MB$), which is used to assist the model in estimating negative impact before to fully model training. A small confidence weight will be assigned to high-negative samples.
@@
  • Extensive experiments are conducted on two manual-annotated datasets and a real-world dataset, showcasing the NPC’s superiority over the state-of-the-art methods. Moreover, with the increasing noise, both quantitative and qualitative results affirm that NPC demonstrates notably higher performance stability compared to previous methods.
@@
Typical image-text matching methods align data from different modalities to measure similarity. Early works (Faghri et al. 2018; Song and Soleymani 2019; Wang et al. 2018; Qian et al. 2021) mainly focus on global alignments. Some prior works (Lee et al. 2018; Li et al. 2019a; Diao et al. 2021; Zhang et al. 2023) adopt attention mechanisms to achieve fine-grained local alignments. Subsequently, many works (Chun et al. 2021; Chun 2023; Li et al. 2022) devote to modeling the many-to-many relationships in image-caption pairs.
Recently, with the success of transformer-based vision and language models (Dosovitskiy et al. 2021; Devlin et al. 2019), vision-language pre-training (VLP) models, such as CLIP (Radford et al. 2021), have shown strong performance in multiple cross-modal tasks (Jiang and Ye 2023; You et al. 2023). Although VLP possesses impressive zero-shot ability, it still reveals vulnerabilities in training with noisy datasets on specific downstream tasks. In this paper, we employ CLIP as our backbone and introduce an anti-noise learning strategy.
@@
Huang et al. (Huang et al. 2021) first tackle noise correspondences, which consider mismatched cross-modal pairs instead of incorrect annotations. Since then, several approaches (Han et al. 2023; Yang et al. 2023; Ye et al. 2022; Ye and Yuen 2020) have developed the noise-rectify process in various cross-modal tasks. They can be categorized into two groups: noise correction and noise re-weighting. Noise correction methods achieve robust learning by correcting the similarity (Han et al. 2023) or correspondence label (Huang et al. 2021; Yang et al. 2023) of noise pairs. The noise re-weighting methods (Qin et al. 2022) degrade the contribution of noise samples to achieve robust learning. All these methods require splitting a noise subset from the original training dataset. Subsequently, they proceed with rectification within this subset. Nonetheless, as noise increases, the imprecise subset division and inaccurate rectification can amplify adverse effects. Different from these works, we sidestep the problem by forecasting per-sample negative impact following the novel NP paradigm.
@@
Fig. 1 provides a clear visual representation, demonstrating that when the noise ratio rises from 20% to 60%, the CLIP’s R@1 performance experiences a steep decline from 82.3% to 66.3%. Therefore, the NPC approach is introduced to enhance the stability and robustness of pre-trained models in tackling noise challenges. The training pipeline, depicted in Fig. 2, comprises the two main components that will be elaborated upon in the subsequent section.
@@
The proposed NPC is evaluated on three benchmark datasets, MSCOCO (Lin et al. 2014), Flickr30K (Young et al. 2014), and CC120K:
@@
  • CC120K. We randomly sample a subset from the real-world dataset Conceptual Captions (Sharma et al. 2018). This dataset is harvested from the Internet, with about 3%-20% incorrect image-text pairs. CC120K contains 120, 851 with a single caption per image. In our experiment, we use 118,851 images for training, 1,000 images for validation, and 1,000 images for testing.
@@
To illustrate the effectiveness, we compare NPC with various approaches, including general cross-modal matching methods SCAN (Lee et al. 2018), SAF (Diao et al. 2021), noise-robust learning methods NCR (Huang et al. 2021), DECL (Qin et al. 2022), BiCro (Yang et al. 2023), and CLIP with fine-tuning (Radford et al. 2021). It is worth noting that CLIP is the baseline of our method. The results are shown in Table 1.
It shows that NPC significantly outperforms all methods across all noise ratios. Notably, on Flickr30K with 60% noise ratio, NPC outperforms the current state-of-the-art approach BiCro with a large R@1 performance gap. To be specific, the R@1 performance of NPC is 15.4% higher than BiCro on image-text matching (i2t), as well as 16.9% higher than BiCro on text-to-image matching (t2i). Compared to the baseline CLIP, NPC has achieved immense improvement in all metrics and benchmarks. Furthermore, as the noise ratio increases, the performance gap between NPC and baseline becomes larger. For instance, on the MSCOCO 1K set, when the noise ratio ranges from 0% to 60%, the R@1 performance gap between NPC and baseline separately increases from 2.3% to 11.2% on i2t, and 3.3% to 13.4% on t2i. This phenomenon is powerful to prove the effectiveness of NPC on robust learning.
@@
Comparison with ViT-B/32 Backbone Methods
Table 2: Comparison with baseline on CC120K.
method image-to-text text-to-image
R@1 R@5 R@10 R@1 R@5 R@10
CLIP 68.8 87.0 92.9 67.8 86.4 90.9
NPC 71.1 92.0 96.2 73.0 90.5 94.8
@@
In Table 2, we compare the NPC with baseline on the CC120K which is with real noisy correspondences. From the results, our proposed method outperforms the baseline by a considerable margin in terms of all metrics. Specifically, NPC is 2.3% and 5.2% higher than CLIP on i2t and t2i R@1, respectively.
[…]
@@ Conclusion
Acknowledgement. This work is partially supported by National Natural Science Foundation of China under Grants (62176188) and the Key Research and Development Program of Hubei Province (2021BAD175)